If sugar cost 0.75$ for 250 ml how much would it cost for 355 ml of sugar ?
Answer by fractalier(6550)

Set up a proportion...
.75 is to 250 ml as x is to 355 ml...OR
.75/250 = x/355
Now cross-multiply and solve for x:
250x = .75(355)
x = .75(355) / 250 = $1.065